# Conceptualization, Methodology, Software, Writing

### v1

@@ 4.5 Qualitative Comparisons
The effectiveness of the proposed residual attention module is further demonstrated through a qualitative evaluation of the features learned by ViT-B and ReViT-B using the MHSA mechanism, both with and without residual attention, respectively. To enable such a comparison, we employ the GradCAM algorithm and apply it to the last two MHSA layers of ViT-B and ReViT-B trained on ImageNet1k with samples from the validation set of the same dataset. It’s worth emphasizing that layers 11 and 12 were selected due to their high globality, aligning with the objective of this experiment, which is to highlight how ReViT-B incorporates low-level features into its learned representations compared to ViT-B. To proceed, the feature maps of both models and the input images used to obtain such features are illustrated in Fig. 8. As can be noticed, the feature maps extracted from ViT-B are poor in details and lack the presence of low-level features like shapes and edges. This is caused by the globalization of its attention mechanism, which provokes the feature collapsing phenomenon. This also proves the observation made by authors in [8], which states that ViT-based architectures relying on the standard MHSA often exhibit limited attention variability, resulting in reduced feature diversity or even feature collapse [32]. On the contrary, ReViT-B, equipped with residual attention, displays greater feature diversity with feature maps almost completely aligned with the entire region where the object of interest is located, preserving the same shape. This phenomenon suggests that ReViT-B can incorporate low-level information like shapes and edges with the global context of the scene. Additionally, such ability can be attributed to residual attention. Thanks to this mechanism, the network can propagate attention maps from shallow layers to deeper layers, slowing down the globalization of attention, improving feature diversity, and incorporating low-level characteristics into the global context. Notably, this augmented feature diversity aligns with ReViT’s improved performance in terms of image classification accuracy compared to the standard ViT architecture.
@@ 4.6 Object Detection and Instance Segmentation
To assess the general applicability of the residual attention module, we test its knowledge transfer ability in downstream tasks. As such, further experiments were conducted on spatial-aware networks, i.e., ReSwin-T and ReMViTv2-T, for the tasks of object detection and instance segmentation, achieving increased results. The evaluation was performed on the COCO2017 dataset for both tasks using the COCO average precision (AP) metric, and the input images were resized with $480\leq H\leq 800$ and $W=1333$. For fair comparisons with the solutions proposed in [24] and [20], the Cascade Mask-RCNN [3] and the Mask R-CNN [13] models were used with the ReSwin-T and ReMViTv2-T as backbones, respectively. Additionally, both ReSwin-T and ReMViTv2-T were initialized with the weights learned during ImageNet1K training, hence, knowledge transferring. Table 7 summarizes results on the subset of 5000 validation images. For the object detection task, the ReSwin-T model performance is higher for $AP_{box}$ and $AP_{box}^{50}$ with respect to the original Swin-T architecture. Regarding the ReMViTv2-T model, the performance improves for all the considered APs compared to the original MViTv2-T architecture. For the instance segmentation task, ReSwin-T and ReMViTv2-T with residual attention increase the overall performance in all the metrics compared to their original works, even though by a small margin.
Framework Backbone $AP_{box}$ $AP_{box}^{50}$ $AP_{box}^{75}$ $AP_{mask}$ $AP_{mask}^{50}$ $AP_{mask}^{75}$ #Params
@@ 5 Ablation Study
In this section, we perform an ablation study to determine the significance of balancing the past and current attention information in the proposed residual attention. We will specifically examine the impact of $\alpha$ on various image classification benchmarks including CIFAR10, CIFAR100, Oxford Flowers-102, and Oxford-IIIT Pet.
@@ 5.1 Inspecting the effect of $\alpha$
These experiments encompass an assessment of ReViT’s performance in classification tasks, with the gating variable held constant while manually adjusting its values to observe the model’s behavior. Specifically, ReViT is tested across four classification benchmarks: CIFAR10, CIFAR100, Oxford Flowers-102, and Oxford-IIIT Pet, with $\alpha$ values spanning from 0 to 1. With 0 being the lower extremity considering only past attention and 1 being the other extremity completely ignoring past attention (i.e., ViT). The outcomes of these experiments are graphically presented in Fig. 9. For $\alpha=0$, according to Eq. (10), ReViT relies only on the attention information from the first layer, and as expected, the performance decreases, especially on complex datasets like CIFAR100 and Oxford Flowers-102. Such performance collapse is explained by the fact that for an $\alpha=0$, ReViT lacks global context. For the other values of $\alpha$, notably, in datasets characterized by a modest number of classes and a surplus of examples per class, such as CIFAR10, the impact is relatively subdued. However, in a more demanding dataset featuring significant inter-class and intra-class variations compared to CIFAR10, like Oxford-IIIT Pet, the role of $\alpha$ becomes more pivotal, exemplified by the model’s sensitivity to varying $\alpha$ values affecting its performance. Furthermore, in datasets like CIFAR100 and Oxford Flowers-102, which are even more complex, the effect of $\alpha$ on ReViT becomes even more important and manifests quite similarly with the model attaining peak performance at $\alpha=0.75$, and experiencing a sharp decline as it approaches 1. Considering that an $\alpha$ value close to 1 typically does not yield the optimal performance across all datasets, it becomes evident that past attention information is crucial for learning better representation, and $\alpha$ assumes a vital role within ReViT. It balances the integration of past and current attention information, facilitating the extraction of high-quality image representations encompassing high and low-level features.
@@ 6 Conclusion
In this study, we introduce the innovative Residual Attention Vision Transformer (ReViT) network, which integrates residual attention learning to enhance the extraction of visual features in Vision Transformer (ViT) architectures. The proposed approach effectively transmits and accumulates attention information from queries and keys across successive Multi-Head Self-Attention (MHSA) layers. Such residual connection prevents diminishing low-level visual features. Additionally, it empowers the model to leverage previously extracted features while learning new ones by decelerating the over-globalization of the attention mechanism. The efficacy and robustness of the proposed methods are evaluated on five popular image classification benchmarks, including ImageNet1K, CIFAR-10, CIFAR-100, Oxford Flowers-102, and Oxford-IIIT Pet, as well as the COCO2017 dataset for object detection and instance segmentation. The results demonstrate that residual attention learning elevates the performance of ViT-based neural models in visual recognition tasks. Furthermore, it adeptly preserves and enhances the ability to uncover and incorporate semantic and spatial relationships when integrated into spatial-aware neural networks, facilitating the identification, localization, and segmentation of individual objects within an image. However, since the introduced residual module contrasts the globalization of the attention mechanism, its efficacy is limited when applied in multi-scale architectures with local awareness, providing less improvement compared to its application in single-scale architecture with global attention. Motivated by such limitations, in future works, refining the residual attention module to better suit multi-scale transformer architectures could provide further insights into its effectiveness in such architectures.

### v2

@@ 4.6 Qualitative Comparisons
The effectiveness of the proposed residual attention module is further demonstrated through a qualitative evaluation of the features learned by ViT-B and ReViT-B using the MHSA mechanism, both with and without residual attention, respectively. To enable such a comparison, we employ the GradCAM algorithm and apply it to the last two MHSA layers of ViT-B and ReViT-B trained on ImageNet1k with samples from the validation set of the same dataset. It’s worth emphasizing that layers 11 and 12 were selected due to their high globality, aligning with the objective of this experiment, which is to highlight how ReViT-B incorporates low-level features into its learned representations compared to ViT-B. To proceed, the feature maps of both models and the input images used to obtain such features are illustrated in Fig. 8. As can be noticed, the feature maps extracted from ViT-B are poor in detail and lack low-level features like shapes and edges. This is caused by the globalization of its attention mechanism, which provokes the feature-collapsing phenomenon. This also proves the observation made by authors in [7], which states that ViT-based architectures relying on the standard MHSA often exhibit limited attention variability, resulting in reduced feature diversity or even feature collapse [35]. On the contrary, ReViT-B, equipped with residual attention, displays greater feature diversity with feature maps almost completely aligned with the entire region where the object of interest is located, preserving the same shape. This phenomenon suggests that ReViT-B can incorporate low-level information like shapes and edges with the global context of the scene. Additionally, such ability can be attributed to residual attention. Thanks to this mechanism, the network can propagate attention maps from shallow layers to deeper layers, slowing down the globalization of attention, improving feature diversity, and incorporating low-level characteristics into the global context. Notably, this augmented feature diversity aligns with ReViT’s improved performance in terms of image classification accuracy compared to the standard ViT architecture and its ability to extract important low-level information in the presence of translation invariances.
@@ 4.7 Object Detection and Instance Segmentation
To assess the general applicability of the residual attention module, we test its knowledge transfer ability in downstream tasks. As such, further experiments were conducted on spatial-aware networks, i.e., ReSwin-T and ReMViTv2-T, for the tasks of object detection and instance segmentation, achieving increased results. The evaluation was performed on the COCO2017 dataset for both tasks using the COCO average precision (AP) metric, and the input images were resized with $480\leq H\leq 800$ and $W=1333$. For fair comparisons with the solutions proposed in [26] and [22], the Cascade Mask-RCNN [1] and the Mask R-CNN [13] models were used with the ReSwin-T and ReMViTv2-T as backbones, respectively. Additionally, both ReSwin-T and ReMViTv2-T were initialized with the weights learned during ImageNet1K training, hence, knowledge transferring. Table 7 summarizes results on the subset of 5000 validation images. For the object detection task, the ReSwin-T model performance is higher for $AP_{box}$ and $AP_{box}^{50}$ with respect to the original Swin-T architecture. Regarding the ReMViTv2-T model, the performance improves for all the considered APs compared to the original MViTv2-T architecture. For the instance segmentation task, ReSwin-T and ReMViTv2-T with residual attention increase the overall performance in all the metrics compared to their original works, even though by a small margin. Additionally, even though residual attention boosts the performances of such models, its impact is lower compared to single-scale architectures due to the multi-scale nature of such architectures and the presence of local information in the produced features.
Framework Backbone $AP_{box}$ $AP_{box}^{50}$ $AP_{box}^{75}$ $AP_{mask}$ $AP_{mask}^{50}$ $AP_{mask}^{75}$Params
@@ 5 Ablation Study
In this section, we perform an ablation study to determine the significance of balancing the past and current information in the proposed residual attention and its effect when is specific to each layer. We will specifically examine the impact of $\alpha$ on various image classification benchmarks, including CIFAR10, CIFAR100, Oxford Flowers-102, and Oxford-IIIT Pet.
@@ 5.1 Inspecting the effect of $\alpha$
These experiments encompass an assessment of ReViT’s performance in classification tasks, with the gating variable held constant while manually adjusting its values to observe the model’s behavior. Specifically, ReViT is tested across four classification benchmarks: CIFAR10, CIFAR100, Oxford Flowers-102, and Oxford-IIIT Pet, with $\alpha$ values spanning from 0 to 1. With 0 being the lower extremity considering only past attention and 1 being the other extremity completely ignoring past attention (i.e., ViT). The outcomes of these experiments are graphically presented in Fig. 9. For $\alpha=0$, according to Eq. (10), ReViT relies only on the attention information from the first layer, and as expected, the performance decreases, especially on complex datasets like CIFAR100 and Oxford Flowers-102. Such performance collapse is explained by the fact that for an $\alpha=0$, ReViT lacks global context. For the other values of $\alpha$, notably in datasets characterized by a modest number of classes and a surplus of examples per class, such as CIFAR10, the impact is relatively subdued. However, in a more demanding dataset featuring significant inter-class and intra-class variations compared to CIFAR10, like Oxford-IIIT Pet, the role of $\alpha$ becomes more pivotal, exemplified by the model’s sensitivity to varying $\alpha$ values affecting its performance. Furthermore, in datasets like CIFAR100 and Oxford Flowers-102, which are even more complex, the effect of $\alpha$ on ReViT becomes even more important and manifests quite similarly with the model attaining peak performance at $\alpha=0.75$, and experiencing a sharp decline as it approaches 1. This increased importance of $\alpha$ on complex dataset is related to the higher necessity of low-level features like colors and shapes needed to discriminate between different classes of such dataset characterized by high inter-class similarity and intra-class variances. Finally, considering that an $\alpha$ value close to 1 typically does not yield the optimal performance across all datasets, it becomes evident that past attention information is crucial for learning better representation, and $\alpha$ assumes a vital role within ReViT. It balances the integration of past and current attention information, facilitating the extraction of high-quality image representations encompassing high and low-level features.
5.2 Learning $\alpha$ for each layer
According to Eq. (10), alpha is a learnable weight scalar that tries to balance the attention flow between consecutive blocks. Additionally, $\alpha$ is kept the same across all layers (i.e., the weight of current and past attention is the same at each layer). Such design choice was intended to resemble a traditional residual connection to show that even a simple operation such as a weighted sum can help with feature diversity when applied to the main catalyst: self-attention. As this operation has shown its benefits, we now give each layer the freedom to choose its balance between past and current attention by making $\alpha$ layer-specific. Mathematically, this means that Eq.(10) is slightly changed as following:
$S_{l}=\begin{cases}\frac{Q_{l}\cdot K_{l}^{T}}{\sqrt{d_{l}}},&\text{if}\ l=0\\ \alpha_{l}(\frac{Q_{l}\cdot K_{l}^{T}}{\sqrt{d_{l}}})+(1-\alpha_{l})(S_{l-1}),&\text{otherwise}\end{cases},$ (13)
Notably, as the computational logic remains the same as in Eq. (10), the subscript $l$ on $\alpha$ (i.e., $\alpha_{l}$) indicates that $\alpha$ is bounded to the specific layer. To proceed with analysing its effect, we train ReViT-B with residual attention implemented according to Eq. (13) on CIFAR10, CIFAR100, Oxford Flowers-102, and Oxford-IIIT Pet datasets and report the per-layer $\alpha_{l}$ value and its impact on the overall performance compared to the default configuration in Fig. 10 (a) and (b) respectively.
Figure 10: The effect of learning layer-specific $\alpha$ values. In (a), the $\alpha_{l}$ values for each layer of ReViT-B on the chosen datasets are reported. In (b), the impact of layers-specific $\alpha$ on the accuracy of ReViT-B on the same datasets is reported.
Upon carefully analyzing the values of $\alpha_{l}$ for different layers as depicted in Fig. 10 (a), we observe a trend where $\alpha_{l}$ decreases as information flows deeper into the layers, placing greater emphasis on past attention information. This behavior aligns with the intended design of $\alpha_{l}$, considering how feature diversity evolves as information progresses towards deeper layers. Specifically, based on our analysis from previous sections, we found that the diversity of features is higher in the top layers of the network. Such observation is further supported by the fact that $\alpha_{l}$ is high in such layers, reducing the necessity of using past attention. In contrast, beyond the fifth layer, where feature diversity diminishes significantly, $\alpha_{l}$ drops drastically, highlighting the importance of the attention from previous layers derived from the interactions of tokens with higher feature diversity. Moreover, we also notice that for datasets with higher image quality, such as Oxford Flowers-102 and Oxford-IIIT Pet, $\alpha_{l}$ decreases slower in the first half of the layers compared to CIFAR10 and CIFAR100 which are known to be low-quality. Such a phenomenon might suggest that in higher-quality images, features tend to lose diversity slower in the shallow layers but still collapse similarly in the deeper ones. Finally, in Fig. 10 (b), we compare the performance of ReViT-B on the four datasets with single learnable $\alpha$ and per layer learnable $\alpha_{l}$. From the results, we conclude that there is an improvement in performance from having learnable $\alpha_{l}$, except in the case of the CIFAR10 dataset. Specifically, for CIFAR10, the obtained results are identical to the setting with global $\alpha$; instead, for CIFAR100, Oxford Flowers-102, and Oxford-IIIT Pet, improvements of 0.16, 0.14 and 0.14, respectively, are obtained by using a per-layer $\alpha$. However, it is worth noting that the margin for improvements in these datasets is low, especially on CIFAR10, with accuracy going beyond 98%.
@@ 6 Conclusion
In this work, we have contributed the Reidual attention Vision Transformer (ReViT) to the family of vision transformers, designed for contrasting the phenomenon of feature collapsing. While previous works approach the targeted problem by augmenting feature connectivity [36], inducing convolution-like bias [7] or shaping the transformer architecture to resemble a CNN [26, 22], ReViT offers a more elegant, simple, target focused and effective solution by addressing the over globalization of the attention mechanism through residual attention connections. These connections give ReViT the ability to propagate attention information across layers, allowing it to capture and retain fine-grained details that are crucial for accurate visual recognition. Extensive experiments on various recognition benchmarks, coupled with in-depth analysis of the residual attention mechanism, showcase ReViT’s effectiveness across diverse tasks and its ability to preserve feature diversity in deeper transformer layers.
The practical implications of ReViT’s enhanced feature diversity are far-reaching. In domains such as medical imaging or specialized industrial inspection, where obtaining large labeled datasets is often infeasible, ReViT offers a promising solution. Its ability to learn effectively even in limited data and capture a rich array of features has the potential to unlock the utility of standard transformer architectures even in such scenarios [21]. Another key strength of ReViT lies in its simple yet effective design. The attention residual connection seamlessly integrates into existing ViT blocks, incurring minimal to no computational overhead while yielding substantial performance gains. This architectural simplicity preserves the core fundamentals of the standard transformer architecture, aligning with the future prospective of designing unified software and hardware solutions given the ability of such architectures to generalize across tasks, domains, and modalities [9]. Something that CNN-like transformer architectures [26, 22, 7, 45] cannot do. Finally, while our study primarily focuses on image classification, the implications of ReViT extend beyond this domain. Preliminary experiments on object detection and instance segmentation, using residual attention in spatial-aware multi-scale transformer models such as Swin [26] and MViTv2 [22], hint at its versatility and potential for broader applications in computer vision.
Limitations. While our study successfully demonstrates a correlation between the globality of attention and feature diversity, it is important to note several limitations. Firstly, residual attention is primarily aimed at mitigating the globalization of the attention mechanism, which has proven beneficial in preserving feature diversity and enhancing overall performance on medium-to-small-scale datasets aligned with our goals. However, its generalization to larger and more diverse networks and datasets remains uncertain. Furthermore, while aiming to preserve the simplicity of self-attention for its ability to generalize across different modalities, the residual attention mechanism we designed was specifically evaluated on image inputs. Although effective for image-based modalities, its generalization to other types of data remains uncertain.
Future Works. While this study has offered valuable insights into methods for mitigating feature collapsing, our findings and limitations suggest several promising avenues for future research and development. Firstly, we aim to validate the generalizability of our approach across larger scales, diverse modalities, and more varied tasks to directly address our identified limitations. It would be interesting to apply ReViT on domains such as medical imaging [47] or aerial imaging [19] to observe the effectiveness of ReViT across domains. Secondly, our current residual attention mechanism straightforwardly combats over-globalization by leveraging attention maps derived from diverse feature interactions. In future work, we intend to shift focus towards identifying dominant tokens within sequences that may contribute to feature unification, offering a new perspective on addressing this challenge. Lastly, it would be interesting to investigate how controlling attention globality and feature diversity impacts the explainability of vision transformers.